# Channel Estimation for RIS-Assisted mmWave Systems via Diffusion Models

### v3

@@ I Introduction
Millimeter-wave (mmWave) communication is a key technology for the sixth generation (6G) communication systems to fully exploit the available spectrum [Wang2023]. Nevertheless, the coverage range of mmWave signals is limited due to their rapid attenuation, which makes them susceptible to blockage. To overcome this limitation, reconfigurable intelligent surfaces (RISs) [Wu2019] have been introduced to extend communication coverage and improve transmission rates. However, the effectiveness of the RIS depends on the acquisition of accurate channel state information (CSI). In general, a RIS consists of passive elements without signal processing capabilities, which poses a significant challenge for channel estimation in RIS-assisted systems.
Conventional methods (e.g., linear minimum mean square error, LMMSE) and compressed sensing (CS)-based methods (e.g., the double-structured orthogonal matching pursuit, DS-OMP [Wei2021]) struggle to achieve accurate channel estimation under low signal-to-noise ratio (SNR) conditions. With the rapid development of deep learning (DL), DL-based methods have been increasingly applied in RIS-assisted systems. For instance, a deep residual network (DRN) is used to denoise the results obtained from the least squares (LS) estimator in [Liu2021]. Furthermore, a global attention residual network (GARN) is introduced to fuse multi-channel information and improve the accuracy of channel estimation [Feng2023]. However, these direct-mapping DL methods require network retraining for different SNRs, limiting their generalizability and practicality.
Recently, diffusion models (DMs) [Ho2020] have garnered significant attention in wireless communications due to their strong capability to model complex data distributions [Wu2024, Guan2024, Zhou2025, Fesl2024]. Among their various applications, channel estimation has received particular attention due to its critical role, especially in RIS-assisted scenarios where cascaded and direct links lead to more complex distributions and greater challenges. In [Tong2024], the authors employ received signals to guide the sampling of DMs for channel estimation. Another approach has been proposed in [Zhang2025], which infers the full channel from the partial channel based on conditional DMs. Moreover, in [Liu2022], the diffusion process is conducted in the spectral domain to improve stability. However, in these works, the sampling process typically starts from the maximum step $T$ without adapting to different signal-to-noise ratio (SNR) conditions, which limits their robustness under varying channel environments. Furthermore, the network architectures adopted in previous works fail to strike a balance among performance, parameter size, and computational complexity. Therefore, achieving accurate and efficient channel estimation in RIS-assisted systems remains a challenging problem that requires further exploration.
@@ II-A Channel Model
As illustrated in Fig. 1, we consider the uplink channel estimation for RIS-assisted mmWave systems. Both the base station (BS) and RIS are equipped with uniform planar arrays (UPAs) with half-wavelength element spacing, while the user equipment (UE) has a single antenna. The BS has $M=M_{h}\times M_{v}$ antennas, and the RIS consists of $N=N_{h}\times N_{v}$ passive elements. The UE-BS, UE-RIS, and RIS-BS channels are denoted as $\mathbf{d}\in\mathbb{C}^{M\times 1}$, $\mathbf{h}\in\mathbb{C}^{N\times 1}$, and $\mathbf{G}\in\mathbb{C}^{M\times N}$, respectively. Following [3GPP2018] and [Basar2021], a clustered statistical channel model is adopted. The RIS-BS channel consists of line-of-sight (LOS) and non-line-of-sight (NLOS) components, i.e., $\mathbf{G}=\mathbf{G}_{\text{LOS}}+\mathbf{G}_{\text{NLOS}}$. These two components are represented as
@@ II-B Signal Model
where $\mathbf{H}=\left[\mathbf{d},\ \mathbf{H}_{c}\right]\in\mathbb{C}^{M\times(N+1)}$. Without loss of generality, the pilot symbol is set to $x_{k}=1$ [Feng2023, Tong2024]. After $K$ time slots, the received signal $\mathbf{Y}\in\mathbb{C}^{M\times K}$ is given by
@@
The RIS reflecting matrix is set as the columns of the discrete Fourier transform (DFT) matrix [Zheng2020]. The objective of channel estimation is to recover the unknown channel $\mathbf{H}$ from the received signal $\mathbf{Y}$. By vectorizing (7), we obtain
@@ III-A Training Algorithm of the Proposed Method
The objective of the diffusion model is to learn the distribution of $\mathbf{x}_{0}$, enabling accurate reconstruction of clean pilot responses from noisy observations. Following the DMs framework [Ho2020], minimizing the variational bound leads to a noise-prediction objective, which is given by
@@
where $D_{\text{KL}}(\cdot)$ denotes the Kullback-Leibler (KL) divergence. The third term captures the gap between the true backward transition and its parameterized approximation. By aligning the parameterized transition with the true conditional distribution of the noisy signal, the model can achieve more accurate denoising and signal reconstruction in the RIS-assisted channel scenario. $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ can be derived via Bayes’ rule as in [Ho2020]. With a specific parameterization of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, $L_{t-1}$ can be calculated in a Rao-Blackwellized fashion with closed-form expressions [Ho2020]. To summarize, the neural network’s loss function can be formulated as
@@
By optimizing (13), the network learns to estimate the noise component in $\mathbf{x}_{t}$, enabling progressive reconstruction of the clean signal during sampling. The inherent angular sparsity of RIS-assisted mmWave channels aligns well with the progressive denoising nature of diffusion models, facilitating more effective learning of the channel structure. Therefore, we adopt the angular-domain representation as in [Fesl2024], and construct the training dataset as $\mathcal{X}=\left\{\mathbf{X}_{n}=\operatorname{fft}(\mathbf{H}_{n}\mathbf{P})\right\}_{n=1}^{N_{\mathrm{train}}}$, where $\operatorname{fft}(\cdot)$ denotes the two-dimensional fast Fourier transform (FFT). In summary, the training algorithm is outlined in Algorithm 1111Consistent with [Ho2020] and [Fesl2024], the vectorized form is employed for theoretical analysis, whereas the matrix form is used in practical implementations..
@@ III-B Sampling Algorithm of the Proposed Method
During the sampling process, the proposed method progressively denoises the signal through the sequential transitions $\mathbf{x}_{t}\rightarrow\mathbf{x}_{t-1}\rightarrow\cdots\rightarrow\mathbf{x}_{0}$. This stepwise refinement restores the dominant angular components of the RIS-assisted channel from noisy observations, leading to more accurate channel estimation. According to [Song2021], $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ is approximated by $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0})$, yielding the following sampling process
@@
Given (9) and (10), when $\sqrt{\bar{\alpha}_{t_{s}}}=\frac{1}{\sqrt{1+\sigma^{2}}}$, the KL divergence satisfies $D_{\mathrm{KL}}(q(\mathbf{x}_{t_{s}}|\mathbf{x}_{0})\parallel p(\mathbf{y}|\mathbf{x}_{0}))=0$. This indicates that the forward process yields signals that follow the same distribution as the received signals. Here, $\mathbf{x}_{t_{s}}$ denotes the signal at step $t_{s}$ in the forward process. Therefore, during the sampling process, we can start from an intermediate step $t_{s}$ rather than from step $T$ as done in [Ho2020], which significantly improves the efficiency. The index $t_{s}$ is computed as
@@
which is referred to as the step-matching mechanism [Wu2024]. The starting step depends on the noise level, with larger indices indicating stronger noise contamination and thus requiring more denoising steps. By incorporating (15), the proposed approach can adapt to varying noise conditions without retraining.
@@ III-C Network Architecture Design
3) Encoder: Motivated by the architectural efficiency and denoising effectiveness of DnCNN [Zhang2017], we adapt its design principles to channel estimation and introduce a novel BRC block. Specifically, this block consists of Batch normalization (BN) layers, ReLU activation functions, and Convolutional layers to extract multi-level features efficiently. To further enhance the model’s capability, we integrate the self-attention mechanism. This tailored design not only improves denoising performance but also reduces model complexity, making it well-suited for practical wireless communication scenarios.
@@ IV-A Experiment Setup
Following the hyperparameter configuration in [Ho2020], we set the total diffusion steps to $T=1000$. The variance schedule $\{\beta_{t}\}$ increases linearly from $\beta_{0}=1\times 10^{-4}$ to $\beta_{T}=0.02$. Our dataset generation code is developed based on the open-source MATLAB simulator SimRIS [Basar2021]. For each scenario, the training set consists of 50,000 channel samples, while the validation and test sets each contain 10,000 channel samples. For training, the Adam optimizer is employed with a cosine warm-up learning rate schedule, where the initial learning rate is set to $1\times 10^{-4}$. The model is trained for 800 epochs with a batch size of 256.
@@ IV-B Results and Analysis
In simulations, the proposed method is compared with the following baselines: (i) the LMMSE estimator; (ii) DS_OMP [Wei2021]; (iii) CDRN [Liu2021]; (iv) LCDM(CNN1) [Fesl2024], which is based on denoising diffusion probabilistic models (DDPMs) with a low-complexity CNN-based backbone; (v) LCDM(CNN2) [Fesl2024], an enhanced version with an increased number of convolutional layers and feature channels to ensure a fair comparison; and (vi) GDM [Tong2024], a method based on guided DMs for RIS channel estimation. For the proposed method, we implement three network architectures: CNN2, the U-Net architecture in [Wu2024], and the proposed BRCNet.
Fig. 3 illustrates the NMSE performance under different SNR conditions. As a representative example, Fig. 3LABEL:sub@fig3:d (0 dB) is analyzed in detail. Unlike the direct-mapping approach of CDRN, the proposed method achieves superior performance through progressive denoising. Compared with LCDM(CNN1) and LCDM(CNN2), Proposed(BRCNet) achieves gains of 9.01 dB and 3.01 dB, respectively, demonstrating its superior effectiveness. By introducing the step-matching mechanism, our method better aligns with channel estimation and therefore outperforms GDM. Furthermore, the comparison between Proposed(CNN2) and Proposed(BRCNet) shows that the tailored BRCNet architecture is more effective than a simple CNN under the same parameter budget. Moreover, the proposed BRCNet achieves performance highly comparable to that of U-Net, with a performance gap of only 0.48 dB, while significantly reducing the number of parameters.
We further conduct an ablation study to evaluate the impact of different diffusion step numbers $T$ and the step-matching mechanism. By changing $T$ in the “Proposed(BRCNet)” configuration from 1000 to 200, we obtain the “T=200(BRCNet)” curve. The results show that $T=1000$ outperforms $T=200$ at low SNRs due to its finer noise-injection and denoising granularity, while the two settings yield similar performance at high SNRs, where only a few denoising steps are sufficient. We also remove the step-matching mechanism to obtain the “w/o step-matching” curve. Without step-matching, the sampling process starts from random Gaussian noise and does not utilize the received signal, resulting in SNR-insensitive performance. In contrast, with step-matching, sampling begins from the received signal itself, enabling effective denoising and yielding better SNR-dependent performance that better aligns with the requirements of the channel estimation task.

### v4

@@ I Introduction
Millimeter-wave (mmWave) communication is a key technology for the sixth generation (6G) communication systems to fully exploit the available spectrum [1]. Nevertheless, the coverage range of mmWave signals is limited due to their rapid attenuation, which makes them susceptible to blockage. To overcome this limitation, reconfigurable intelligent surfaces (RISs) [2] have been introduced to extend communication coverage and improve transmission rates. However, the effectiveness of the RIS depends on the acquisition of accurate channel state information (CSI). In general, a RIS consists of passive elements without signal processing capabilities, which poses a significant challenge for channel estimation in RIS-assisted systems.
Conventional methods (e.g., linear minimum mean square error, LMMSE) and compressed sensing (CS)-based methods (e.g., the double-structured orthogonal matching pursuit, DS-OMP [3]) struggle to achieve accurate channel estimation under low signal-to-noise ratio (SNR) conditions. With the rapid development of deep learning (DL), DL-based methods have been increasingly applied in RIS-assisted systems. For instance, a deep residual network (DRN) is used to denoise the results obtained from the least squares (LS) estimator in [4]. Furthermore, a global attention residual network (GARN) is introduced to fuse multi-channel information and improve the accuracy of channel estimation [5]. However, these direct-mapping DL methods require network retraining for different SNRs, limiting their generalizability and practicality.
Recently, diffusion models (DMs) [6] have garnered significant attention in wireless communications due to their strong capability to model complex data distributions [7, 8, 9, 10]. Among their various applications, channel estimation has received particular attention due to its critical role, especially in RIS-assisted scenarios where cascaded and direct links lead to more complex distributions and greater challenges. In [11], the authors employ received signals to guide the sampling of DMs for channel estimation. Another approach has been proposed in [12], which infers the full channel from the partial channel based on conditional DMs. Moreover, in [13], the diffusion process is conducted in the spectral domain to improve stability. However, in these works, the sampling process typically starts from the maximum step $T$ without adapting to different SNR conditions, which limits their robustness under varying channel environments. Furthermore, the network architectures adopted in previous works fail to strike a balance among performance, parameter size, and computational complexity. Therefore, achieving accurate and efficient channel estimation in RIS-assisted systems remains a challenging problem that requires further exploration.
@@ II-A Channel Model
As illustrated in Fig. 1, we consider the uplink channel estimation for RIS-assisted mmWave systems. Both the base station (BS) and RIS are equipped with uniform planar arrays (UPAs) with half-wavelength element spacing, while the user equipment (UE) has a single antenna. The BS has $M=M_{h}\times M_{v}$ antennas, and the RIS consists of $N=N_{h}\times N_{v}$ passive elements. The UE-BS, UE-RIS, and RIS-BS channels are denoted as $\mathbf{d}\in\mathbb{C}^{M\times 1}$, $\mathbf{h}\in\mathbb{C}^{N\times 1}$, and $\mathbf{G}\in\mathbb{C}^{M\times N}$, respectively. Following [14] and [15], a clustered statistical channel model is adopted. The RIS-BS channel consists of line-of-sight (LOS) and non-line-of-sight (NLOS) components, i.e., $\mathbf{G}=\mathbf{G}_{\text{LOS}}+\mathbf{G}_{\text{NLOS}}$. These two components are represented as
@@ II-B Signal Model
where $\mathbf{H}=\left[\mathbf{d},\ \mathbf{H}_{c}\right]\in\mathbb{C}^{M\times(N+1)}$. Without loss of generality, the pilot symbol is set to $x_{k}=1$ [5, 11]. After $K$ time slots, the received signal $\mathbf{Y}\in\mathbb{C}^{M\times K}$ is given by
@@
The RIS reflecting matrix is set as the columns of the discrete Fourier transform (DFT) matrix [16]. The objective of channel estimation is to recover the unknown channel $\mathbf{H}$ from the received signal $\mathbf{Y}$. By vectorizing (7), we obtain
@@ III-A Training Algorithm of the Proposed Method
The objective of the diffusion model is to learn the distribution of $\mathbf{x}_{0}$, enabling accurate reconstruction of clean pilot responses from noisy observations. Following the DMs framework [6], minimizing the variational bound leads to a noise-prediction objective, which is given by
@@
where $D_{\text{KL}}(\cdot)$ denotes the Kullback-Leibler (KL) divergence. The third term captures the gap between the true backward transition and its parameterized approximation. By aligning the parameterized transition with the true conditional distribution of the noisy signal, the model can achieve more accurate denoising and signal reconstruction in the RIS-assisted channel scenario. $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ can be derived via Bayes’ rule as in [6]. With a specific parameterization of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, $L_{t-1}$ can be calculated in a Rao-Blackwellized fashion with closed-form expressions [6]. To summarize, the neural network’s loss function can be formulated as
@@
By optimizing (13), the network learns to estimate the noise component in $\mathbf{x}_{t}$, enabling progressive reconstruction of the clean signal during sampling. The inherent angular sparsity of RIS-assisted mmWave channels aligns well with the progressive denoising nature of diffusion models, facilitating more effective learning of the channel structure. Therefore, we adopt the angular-domain representation as in [10], and construct the training dataset as $\mathcal{X}=\left\{\mathbf{X}_{n}=\operatorname{fft}(\mathbf{H}_{n}\mathbf{P})\right\}_{n=1}^{N_{\mathrm{train}}}$, where $\operatorname{fft}(\cdot)$ denotes the two-dimensional fast Fourier transform (FFT). In summary, the training algorithm is outlined in Algorithm 1111Consistent with [6] and [10], the vectorized form is employed for theoretical analysis, whereas the matrix form is used in practical implementations..
@@ III-B Sampling Algorithm of the Proposed Method
During the sampling process, the proposed method progressively denoises the signal through the sequential transitions $\mathbf{x}_{t}\rightarrow\mathbf{x}_{t-1}\rightarrow\cdots\rightarrow\mathbf{x}_{0}$. This stepwise refinement restores the dominant angular components of the RIS-assisted channel from noisy observations, leading to more accurate channel estimation. According to [17], $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ is approximated by $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0})$, yielding the following sampling process
@@
Given (9) and (10), when $\sqrt{\bar{\alpha}_{t_{s}}}=\frac{1}{\sqrt{1+\sigma^{2}}}$, the KL divergence satisfies $D_{\mathrm{KL}}(q(\mathbf{x}_{t_{s}}|\mathbf{x}_{0})\parallel p(\mathbf{y}|\mathbf{x}_{0}))=0$. This indicates that the forward process yields signals that follow the same distribution as the received signals. Here, $\mathbf{x}_{t_{s}}$ denotes the signal at step $t_{s}$ in the forward process. Therefore, during the sampling process, we can start from an intermediate step $t_{s}$ rather than from step $T$ as done in [6], which significantly improves the efficiency. The index $t_{s}$ is computed as
@@
which is referred to as the step-matching mechanism [7]. The starting step depends on the noise level, with larger indices indicating stronger noise contamination and thus requiring more denoising steps. By incorporating (15), the proposed approach can adapt to varying noise conditions without retraining.
@@ III-C Network Architecture Design
3) Encoder: Motivated by the architectural efficiency and denoising effectiveness of DnCNN [18], we adapt its design principles to channel estimation and introduce a novel BRC block. Specifically, this block consists of Batch normalization (BN) layers, ReLU activation functions, and Convolutional layers to extract multi-level features efficiently. To further enhance the model’s capability, we integrate the self-attention mechanism. This tailored design not only improves denoising performance but also reduces model complexity, making it well-suited for practical wireless communication scenarios.
@@ IV-A Experiment Setup
Following the configuration in [6], we set the total diffusion steps to $T=1000$. The variance schedule $\{\beta_{t}\}$ increases linearly from $\beta_{0}=1\times 10^{-4}$ to $\beta_{T}=0.02$. Our dataset generation code is developed based on the open-source MATLAB simulator SimRIS [15]. For each scenario, the training set consists of 50,000 channel samples, while the validation and test sets each contain 10,000 channel samples. For training, the Adam optimizer is employed with a cosine warm-up learning rate schedule, where the initial learning rate is set to $1\times 10^{-4}$. The model is trained for 800 epochs with a batch size of 256.
@@ IV-B Results and Analysis
In simulations, the proposed method is compared with the following baselines: (i) the LMMSE estimator; (ii) DS-OMP [3]; (iii) CDRN [4]; (iv) LCDM (CNN1) [10], which is based on denoising diffusion probabilistic models (DDPMs) with a low-complexity CNN-based backbone; (v) LCDM (CNN2) [10], an enhanced version with an increased number of convolutional layers and feature channels to ensure a fair comparison; and (vi) GDM [11], a method based on guided DMs for RIS channel estimation. For the proposed method, we implement three network architectures: CNN2, the U-Net architecture in [7], and the proposed BRCNet.
Fig. 3 illustrates the NMSE performance under different SNR conditions. As an example, Fig. 3LABEL:sub@fig3:d (0 dB) is analyzed in detail. Unlike the direct-mapping approach of CDRN, the proposed method achieves superior performance through progressive denoising. Compared with LCDM (CNN1) and LCDM (CNN2), Proposed (BRCNet) achieves gains of 9.01 dB and 3.01 dB, respectively, demonstrating its superior effectiveness. By introducing the step-matching mechanism, our method better aligns with channel estimation and therefore outperforms GDM. Furthermore, the comparison between Proposed (CNN2) and Proposed (BRCNet) shows that the tailored BRCNet architecture is more effective than a simple CNN under the same parameter budget. Moreover, the proposed BRCNet achieves performance highly comparable to that of U-Net, with a performance gap of only 0.48 dB, while significantly reducing the number of parameters.
We further conduct an ablation study to evaluate the impact of different diffusion step numbers $T$ and the step-matching mechanism. By changing $T$ in the “Proposed (BRCNet)” from 1000 to 200, we obtain the “T=200 (BRCNet)” curve. The results show that $T=1000$ outperforms $T=200$ at low SNRs due to its finer noise-injection and denoising granularity, while the two settings yield similar performance at high SNRs, where only a few denoising steps are sufficient. We also remove the step-matching mechanism to obtain the “w/o step-matching” curve. Without step-matching, the sampling process starts from random Gaussian noise and does not utilize the received signal, resulting in SNR-insensitive performance. In contrast, with step-matching, sampling begins from the received signal, enabling effective denoising and yielding better performance that better aligns with the requirements of the channel estimation.